Saniaheady21:

A table contains the ordered pairs (3, 42.25) and (5, 50.75). If the relationship in the table is linear, explain how to find the initial value.

surjithayer:

\[slope=\frac{ 50.75-42.25 }{ 5-3}=\frac{ 8.5 }{ 2 }=4.25\] eq. of line through (3,42.25) is y-42.25=4.25(x-3) when x=0 y-42.25=-12.75 y=42.25-12.75=29.5 initial value=?